### v2

@@ 2.3 Contribution of assimilating different observations
As illustrated in Figure S1, a pivotal innovation of XiChen lies in its scalable capability to integrate multi-modal observational data within a cascade DA framework for the precise estimation of initial conditions. With the continuous expansion in both the number and diversity of observational modalities, a pressing question emerges: which specific observations exert the most significant influence on medium-range weather forecasting accuracy [laloyaux2025using]? To address this inquiry, we designed an ablation experiment aimed at quantifying the marginal impact of omitting individual observation sources on 10-day medium-range forecast errors.
@@
Figure 3: Headline scorecards of the average normalized RMSEs for 10-day medium-range weather forecasting, based on the assimilation of different observations. The colors denote the percentage difference relative to the baseline, which assimilates all observations, including GDAS prepbufr, AMSU-A, MHS, ASCAT, and SATWND. Blue shading indicates a reduction in normalized RMSE, while red shading indicates an increase.
@@
Similarly, the exclusion of data from channels 3 to 5 of the MHS, which are predominantly utilized for observing atmospheric humidity profiles [zou2017impacts], led to increased errors in upper-air specific humidity. This finding underscores the effectiveness of the observation operator within the XiChen framework in accurately encapsulating MHS’s sensitivity to atmospheric humidity profiles. Moreover, the omission of MHS observations was observed to introduce elevated errors across other variables to a certain degree. This observation further emphasizes XiChen’s capability to leverage its learned equilibrium constraints, effectively redistributing the specific humidity adjustments to other related variables within the system.
@@
Figure 4: Headline scorecards of the global-averaged latitude-weighted RMSEs for 10-day medium-range weather forecasting, based on the assimilation of different observations. The 10-day medium-range forecasts based on analysis fields that assimilate all observational data—including GDAS prepbufr, AMSU-A, MHS, ASCAT, and SATWND—are compared with the forecasts based on analysis fields that assimilate only the GDAS prepbufr. The colors indicate the percentage difference relative to the baseline, which assimilates only the GDAS prepbufr. Blue shading represents a reduction in RMSE, while red shading represents an increase.
@@
Notably, the exclusion of all satellite observations results in a substantially greater increase in forecast errors compared to the removal of a single satellite observation type. This highlights the robust capability of XiChen to effectively integrate a diverse array of satellite observations, thereby enhancing forecast accuracy. Additionally, Figure 4 illustrates the influence of additional assimilating data from MHS, ASCAT, AMSU-A, and SATWND, alongside the GDAS prepbufr, on the RMSE of 10-day medium-range weather forecasts. Across nearly the entire forecast horizon, the incorporation of satellite observations markedly reduces forecast errors for all variables. This reduction is particularly pronounced for upper-atmospheric variables, such as geopotential height, temperature, and specific humidity, underscoring the critical role of satellite data in improving the accuracy of medium-range weather predictions.
2.4 Physical Consistency of the Assimilation
This section examines how XiChen account for the impacts of assimilating a single perturbation observation on the background field and evaluates whether this process adheres to fundamental physical principles.
In this experiment, two separate DA runs were conducted. The first run used ERA5 as the background field to assimilate raw satellite observations. In the second run, a perturbation is introduced to the raw satellite inputs across all observation channels individually. The differences between the two DA results illustrate the impact of the perturbation on the analysis fields. The DA time was set to 00 UTC on July 24, 2023, and perturbation experiments were conducted separately for AMSU-A and MHS. The perturbation location, selected at 20 N, 120 E, was near the path of Typhoon Doksuri during this period. The perturbation magnitude was set to +5 K, and independent experiments were performed for each satellite channel.
Figures 5 and 6 present the horizontal and vertical distributions of analysis field variations resulting from individually perturbing three AMSU-A channels (5, 6, and 7) and three MHS channels (3, 4, and 5), respectively. For AMSU-A, increases in brightness temperature correspond to rising atmospheric temperatures, consistent with radiative transfer theory, demonstrating that XiChen has effectively learned AMSU-A’s temperature sensitivity. For MHS, increases in brightness temperature correspond to decreases in specific humidity, indicating that XiChen has similarly captured MHS’s sensitivity to atmospheric moisture. In both figures, the analysis increments exhibit propagating variations horizontally and vertically and display localized, anisotropic structures around the observations that align with the background geopotential field, reflecting the flow-dependent nature of the XiChen system. Additionally, for AMSU-A, anticyclonic wind responses surrounding temperature increments imply that XiChen adheres to certain equilibrium constraints. Overall, these results are consistent with findings from 4DEnVar [zhu2022four].
Figure 5: Changes in the analysis fields resulting from a 5 K perturbation to the AMSU-A observation at the selected location, where the background field is the ERA5 data at 00 UTC on July 24, 2023. The perturbation was introduced at 03 UTC, three hours after the analysis time, near Typhoon Doksuri at 20 N, and 120 E (indicated by the yellow dot). The first row displays the horizontal spatial distribution of analysis changes for channels 5 to 7 at 700, 200, and 200 hPa. Shading represents temperature differences (units: $K$), while vectors indicate wind differences (units: $ms^{-1}$) resulting from the assimilation of single-point perturbation observations. The solid contour illustrates the background geopotential field (units: $m^{2}s^{-2}$). The second row displays the vertical distribution along the same west-east cross-section.
Figure 6: Changes in the analysis fields resulting from a 5 K perturbation to the MHS observation at the selected location, where the background field is the ERA5 data at 00 UTC on July 24, 2023. The perturbation was introduced at 03 UTC, three hours after the analysis time, near Typhoon Doksuri at 20 N, and 120 E (indicated by the yellow dot). The first row shows the horizontal spatial distribution of analysis changes for channels 3 to 5 at 300, 500, and 500 hPa. Shading represents differences in specific humidity (units: $gkg^{-1}$), while vectors indicate wind differences (units: $ms^{-1}$) resulting from the assimilation of single-point perturbation observations. The solid contour illustrates the background geopotential field (units: $m^{2}s^{-2}$). The second row displays the vertical distribution along the same west-east cross-section.
2.5 Contribution of the 4DVar knowledge
In this section, we investigate the role of 4DVar knowledge and examine the impact of the trained XiChen model on DA and forecasting performance under different DAWs. In our experiments, the DAW is defined as left-closed and right-open. For instance, a 3-hour DAW incorporates observations only at the initial time, while a 12-hour DAW incorporates observations at the initial time and three additional observations at 3-hour intervals thereafter. We conducted experiments with 3-hour DAW (observations at [0h]), 6-hour DAW (observations at [0h, 3h]), 9-hour DAW(observations at [0h, 3h, 6h]), and 12-hour DAW (observations at [0h, 3h. 6h. 9h]) to evaluate the effect of incorporating 4DVar knowledge over varying periods using the 4DVar cost function gradients.
Figure 7: Comparison of the average latitude RMSE (first and third rows) and Bias (second and fourth rows) of the one-year DA cycle using 3-hourly DAW (blue lines), 6-houly DAW (green lines), 9-hourly DAW (yellow lines), and 12-hourly DAW (red lines) for 4 surface variables, such as MSLP, T2M, U10, and V10, and 4 upper-air variables, including Z500, T850, T500, Z850, and Q500, using testing data during 2023. All analysis fields are evaluated against the ERA5 reanalysis dataset.
We performed a one-year DA cycle experiment, with results presented in Figure 7. The findings indicate that as the DAW length increases, the analysis error decreases, highlighting the positive impact of including more temporal observations on XiChen’s performance. Notably, even without the 4DVar knowledge constraint (DAW [0h]), XiChen maintained a stable DA cycle over one year, despite producing larger analysis errors. This result further demonstrates the robustness of the XiChen model. Additionally, we conducted a 10-day medium-range weather forecast experiment, with results provided in Figure S8. Forecasts generated from analysis fields with longer DAWs exhibited lower RMSE and higher ACC, further confirming that incorporating longer temporal information through 4DVar knowledge enhances forecasting accuracy. These results collectively demonstrate the effective improvement in analysis and forecasting performance achieved by introducing 4DVar knowledge.
@@ 2.6 Assimilation and forecasting of tropical cyclones
Figure 8: Assimilation and Forecasting of Tropical Cyclones. (a) Tropical cyclone (TC) trajectory tracks from IBTrACS ground truth data for the entire year of 2023. (b) TC trajectory tracks (blue lines) from the ERA5 reanalysis for 2023. (c) TC trajectory tracks (blue lines) from the XiChen analysis fields obtained by assimilating GDAS prepbufr, AMSU-A, and MHS during a one-year data assimilation (DA) cycle. (d) Comparison of XiChen performance in terms of mean direct position error over 5-day TC forecasts in 2023. (e) Comparison of XiChen performance in terms of average maximum wind speed error over 5-day TC forecasts in 2023. (f) Comparison of XiChen performance in terms of average minimum MSLP error over 5-day TC forecasts in 2023. In panels (d) through (f), the blue lines denote the outcomes initialized with ERA5 data, whereas the yellow and red lines correspond to results derived from analysis fields obtained by assimilating GDAS prepbufr exclusively and by further incorporating all satellite observations, respectively.
@@
To assess XiChen’s ability to assimilate and forecast TC trajectories, we employ the TempestExtremes tracker [ullrich2021tempestextremes] to identify TC tracks. For comparison with the IBTrACS dataset and ERA5 reanalysis, we adopt the configuration used in NeuralGCM [kochkov2024neural], with parameter details listed in Table S9. To maintain consistency with IBTrACS, we exclude any TCs identified by this configuration that are not present in the IBTrACS dataset, ensuring that only IBTrACS TCs are included in our results. For details, please see the Supplementary Text.
@@
This study evaluates XiChen’s performance in assimilating and forecasting TC trajectories. Figures 8(a)-(c) provide a comparative analysis of TC tracks from different datasets and XiChen’s analysis fields. Figure 8(a) depicts the TC trajectories from the IBTrACS ground truth dataset for 2023, including 81 TCs. The blue lines in Figure 8(b) depict the TC trajectories identified in the ERA5 reanalysis dataset for 2023 that correspond with TC trajectories in IBTrACS, comprising a total of 52 tropical cyclones. The ERA5 dataset successfully identified 64.2% of the TCs documented in IBTrACS, with an average trajectory deviation of 0.66 degrees. The blue lines in Figure 8(c) represent TC trajectories generated by XiChen during the one-year DA cycle. The XiChen analysis fields successfully identified 61.73% of the TCs documented in IBTrACS, with an average trajectory deviation of 0.79 degrees. These results demonstrate that XiChen effectively leverages multi-source observational data to retain TC information within the analysis field. Notably, the TC trajectories generated by XiChen include some TCs not identified in ERA5, suggesting that XiChen not only reproduces ERA5 data but also exhibits a degree of generalization capability. Figure S9 displays the results of TC trajectory identification from forecasting fields across different forecast lead times.
To quantitatively evaluate the impact of satellite observations on track and intensity errors in XiChen‘s TC forecasts, we conducted 5-day predictions starting from the onset times of all typhoons in the IBTrACS dataset. Tracking and error calculations were performed at 6-hour intervals. The typhoon tracking methodology follows the scheme established by the Pangu-Weather system [bi2023accurate]. Figures 8(d)–(f) compare forecasts initiated with three different initial fields: (1) ERA5 data (blue), (2) assimilation of only GDAS prepbufr data (yellow), and (3) simultaneous assimilation of GDAS prepbufr, MHS, ASCAT, AMSU-A, and SATWND data (red). Figure 8(d) shows that incorporating satellite observations into GDAS prepbufr assimilation significantly improves the accuracy of TC track forecasts. However, since the forecast model was trained using ERA5 data, its performance when using ERA5 as the initial field remains superior. In Figures 8(e) and 8(f), XiChen demonstrates similar maximum wind speed errors and minimum mean sea level pressure errors across all three initial fields. This is primarily due to biases in the ERA5 dataset regarding TC intensity, which were learned by XiChen during training.
@@ 3 Discussion
The development of XiChen represents a significant breakthrough in fully AI-driven global weather forecasting. Leveraging the scalability of foundation models, XiChen is constructed by a pretrained medium-range forecasting model and fine-tuned to serve as observation operators and DA models. The entire pipeline of XiChen, from DA to 10-day weather forecasting, can be accomplished within only 15 seconds, representing an exceeding 400-fold acceleration compared to the operational NWP system [bauer2020ecmwf]. By integrating GDAS prepbufr, MHS, ASCAT, AMSU-A, and SATWND observations, XiChen offers data-driven global weather forecasting capabilities and demonstrates superior performance compared to the GFS in both DA and forecasting tasks for several key variables. Notably, XiChen achieves a skillful lead time of over 8.75 days, which surpasses that of the GFS. Moreover, XiChen exhibits high scalability: when new satellite observations are introduced, it requires only fine-tuning the foundation model to the respective observation operator, followed by fine-tuning the corresponding assimilation model using the cascade DA framework proposed in this study. Similarly, if an observation becomes temporarily unavailable, the assimilation process can exclude the corresponding model from the pipeline. These capabilities are critical for advancing AI-based DA and forecasting toward practical operationalization. Finally, we verified that XiChen assimilation incorporates physical equilibrium constraints and flow-dependent characteristics through single-point perturbation DA experiments, further confirming its reliability. This achievement underscores the transformative potential of AI in advancing operational weather forecasting and highlights the necessity of sustained investments in AI research to tackle the complexities of Earth system modeling. As weather forecasting enters a new era, XiChen establishes a robust foundation for applying AI to deepen our forecasting of planetary systems, ultimately guiding decisions with profound implications for society and the environment. However, this does not imply that AI-driven global weather forecasting systems can fully replace NWP. Most AI-driven weather forecasting systems, including XiChen, rely heavily on reanalysis fields produced by NWP systems for training. Consequently, fully AI-driven global weather forecasting, exemplified by XiChen, serves as a valuable complement to NWP rather than a replacement.
@@
Although XiChen demonstrates significant potential for practical applications, it also exhibits certain limitations. First, the model experiences smoothing effects during both the DA and weather forecasting processes, a common challenge in AI-based weather forecasting models [bonavita2024some]. Addressing this issue requires further dedicated research. Potential solutions include designing more suitable loss functions [xu2024extremecast], incorporating physical constraints [xu2024weathergft], or employing post-processing techniques based on generative models [zhong2024fuxiextreme]. Second, the current version of XiChen has been trained and tested exclusively at a resolution of 1.40625. To predict finer-scale weather phenomena, future work should focus on extending this resolution to higher levels, such as 0.25 [ailabchen2025fengwu] or even 0.09 [han2024fengwughr]. Third, the current implementation of XiChen integrates satellite observations from only four satellite instruments—MHS, ASCAT, AMSU-A, and SATWND—significantly fewer than those utilized in operational NWP systems [zhong2024fuxiextreme]. In the future, incorporating additional indirect observational data could significantly enhance the performance of the analysis field. Finally, the model produces only deterministic weather forecasts. However, making critical decisions requires awareness of all possible weather scenarios and their associated probabilities. Therefore, it is necessary to extend the model to a probabilistic version. Addressing this limitation in the future may involve reconstructing the model to adopt a probabilistic framework [price2025probabilistic] or incorporating perturbations through the deterministic XiChen system [chen2024machine]. Ultimately, our goal is to develop a fully AI-driven global weather forecasting system that rivals the IFS HRES model in both assimilation and forecasting capabilities. Achieving this ambitious goal necessitates expertise in atmospheric science and the seamless integration of strengths from the atmospheric and AI communities to work together.
4 Medthod

### v3

@@ 2.3 The impact of assimilating different observations on medium-range global weather forecasts
A pivotal innovation of XiChen lies in its scalable and flexible capability to integrate heterogeneous observations within a cascade DA framework for the accurate estimation of initial conditions (see Figure S1 for details). With the continuous expansion in both the number and diversity of observational modalities, a pressing question emerges: which specific observations exert the most significant influence on medium-range global weather forecasting accuracy [laloyaux2025using]. To answer this question, we designed an ablation experiment aimed at quantifying the marginal impact of omitting individual observation sources on 10-day medium-range forecast errors.
@@
Figure 4: Headline scorecards of the average normalized RMSEs for 10-day medium-range global weather forecasting, based on the assimilation of different observations. The colors denote the percentage difference relative to the baseline, which assimilates all observations, including GDAS prepbufr, AMSU-A, MHS, ASCAT, and SATWND. Blue shading indicates a reduction in normalized RMSE, while red shading indicates an increase.
@@
Similarly, the exclusion of data from channels 3 to 5 of the MHS, which are predominantly utilized for observing atmospheric humidity profiles [zou2017impacts], led to increased errors in upper-air specific humidity. This demonstrates the effectiveness of the observation operator within the XiChen framework in accurately encapsulating MHS’s sensitivity to atmospheric humidity profiles. Moreover, the omission of MHS observations was observed to introduce elevated errors across other variables to a certain degree. This observation further emphasizes XiChen’s capability to leverage its learned equilibrium constraints, effectively redistributing the specific humidity adjustments to other related variables within the system.
@@
Figure 5: Headline scorecards of the global-averaged latitude-weighted RMSEs for 10-day medium-range global weather forecasting, based on the assimilation of different observations. The 10-day medium-range forecasts based on analysis fields that assimilate all observations are compared with the forecasts based on analysis fields that assimilate only the GDAS prepbufr. The colors indicate the percentage difference relative to the baseline, which assimilates only the GDAS prepbufr. Blue shading represents a reduction in RMSE, while red shading represents an increase.
@@
Notably, the exclusion of all satellite observations results in a substantially greater increase in forecast errors compared to the removal of a single satellite observation type. This highlights the robust capability of XiChen to effectively integrate a diverse array of satellite observations, thereby enhancing forecast accuracy. Additionally, Figure 5 illustrates the influence of additionally assimilating data from MHS, ASCAT, AMSU-A, and SATWND, alongside the GDAS prepbufr, on the RMSE of 10-day medium-range global weather forecasts. Across nearly the entire forecast horizon, the incorporation of satellite observations markedly reduces forecast errors for all variables. This reduction is particularly pronounced for upper-atmospheric variables, such as geopotential, temperature, and specific humidity, underscoring the critical role of satellite data in improving the accuracy of medium-range global weather predictions.
@@ 2.4 Assimilation and forecasting of tropical cyclones
To assess XiChen’s ability to assimilate and forecast TC trajectories, we employ the TempestExtremes tracker [ullrich2021tempestextremes] to identify TC tracks. For comparison with the IBTrACS [Kenneth2010IBTrACS] dataset and ERA5 reanalysis, we adopt the configuration used in NeuralGCM [kochkov2024neural], with parameter details listed in Table S9. To maintain consistency with IBTrACS, we exclude any TCs identified by this configuration that are not present in the IBTrACS dataset, ensuring that only IBTrACS TCs are included in our results (see the Supplementary Text for details).
Figure 6: Assimilation and Forecasting of TCs. (a) TC trajectory tracks from IBTrACS ground truth data for the entire year of 2023. (b) TC trajectory tracks (blue lines) from the ERA5 reanalysis for 2023. (c) TC trajectory tracks (blue lines) from the XiChen analysis fields obtained by assimilating all observations during a one-year DA cycle. (d) Comparison of XiChen’s performance in terms of mean direct position error over 5-day TC forecasts in 2023. (e) Comparison of XiChen’s performance in terms of average maximum wind speed error over 5-day TC forecasts in 2023. (f) Comparison of XiChen’s performance in terms of average minimum MSLP error over 5-day TC forecasts in 2023. In panels (d) through (f), the blue lines denote the outcomes initialized with ERA5 data, whereas the yellow and red lines correspond to results derived from analysis fields obtained by assimilating GDAS prepbufr exclusively and by further incorporating all satellite observations, respectively.
@@
Figures 6(a)-(c) provide a comparative analysis of TC tracks from different datasets and XiChen’s analysis fields. Figure 6(a) depicts the TC trajectories from the IBTrACS ground truth dataset for 2023, including 81 TCs. The blue lines in Figure 6(b) depict the TC trajectories identified in the ERA5 reanalysis dataset for 2023 that correspond with TC trajectories in IBTrACS, comprising a total of 52 tropical cyclones. The ERA5 dataset successfully identified 64.2% of the TCs documented in IBTrACS, with an average trajectory deviation of 0.66 degrees. The blue lines in Figure 6(c) represent TC trajectories generated by XiChen during the one-year DA cycle. The XiChen analysis fields successfully identified 61.73% of the TCs documented in IBTrACS, with an average trajectory deviation of 0.79 degrees. These results demonstrate that XiChen effectively leverages multi-source observations to retain TC information within the analysis field. Notably, the TC trajectories generated by XiChen include some TCs not identified in ERA5, suggesting that XiChen not only reproduces ERA5 data but also exhibits a degree of generalization capability. As lead time increases, the number of TC tracks identified by XiChen decreases to a certain extent. This is primarily due to the smoothing effect inherent in ML-based weather forecasting (see Figure S11 for details).
To quantitatively evaluate the impact of satellite observations on track and intensity errors in XiChen’s TC forecasts, we conducted 5-day predictions starting from the onset times of all typhoons in the IBTrACS dataset. Tracking and error calculations were performed at 6-hour intervals. The typhoon tracking methodology follows the scheme established by the Pangu-Weather system [bi2023accurate]. The sensitivity of tropical cyclone forecasts to different initial fields is examined using three configurations: ERA5 data, assimilation of GDAS prepbufr observations alone, and joint assimilation of GDAS prepbufr with all satellite observations (Figures 6(d)–(f)). The inclusion of satellite observations leads to a noticeable reduction in TC trajectories forecast errors relative to using GDAS prepbufr alone (Figure 6(d)), suggesting that satellite data provide additional dynamical constraints beneficial for TC trajectory prediction. Forecasts initialized from ERA5 nonetheless exhibit the highest overall skill, which is likely related to the fact that the weather forecasting model was trained solely on ERA5-based states. For intensity-related metrics, including maximum wind speed and minimum MSLP, differences among the three initialization strategies are relatively small (Figures 6(e) and (f)). One possible explanation is that systematic biases in the TC intensity estimation of ERA5 data are inherited by XiChen during training, thereby limiting its sensitivity to alternative initial fields. In addition, many existing ML models tend to produce overly smoothed representations of atmospheric states, which can further hinder the accurate prediction of extreme intensity values [lam2023learning, fudanchen2023fuxi]. Overcoming these limitations remains an open problem for future work.
@@ 3 Discussion
Advances in weather forecasting rely on the integration of new observations, and the rapid development of ML with the increasing availability of atmospheric observations offers a new opportunity to revolutionize weather forecasting. In this study, we construct XiChen, a global weather data-to-forecast ML system via 4DVar gradient-guided observation-flexible assimilation. We demonstrate that the gradient of the 4DVar cost function provides a unified information interface for diverse heterogeneous observations, which enables physically consistent and observation-flexible DA for global weather forecasting. We verified that XiChen’s assimilation incorporates physical equilibrium constraints and flow-dependent characteristics through single-point perturbation DA experiments. By assimilating GDAS prepbufr, MHS, ASCAT, AMSU-A, and SATWND observations, XiChen offers global weather observation-to-forecast capabilities and demonstrates superior performance compared to the GFS in both DA and forecasting tasks for several key variables. Notably, XiChen achieves a skillful lead time of over 8.75 days, which surpasses that of the GFS. Beyond weather forecasting performance, this work illustrates how physics-informed ML can serve as a bridge between heterogeneous Earth observations and data-driven models, offering a scalable pathway toward next-generation Earth system prediction.
Moreover, XiChen demonstrates exceptional flexibility by streamlining the assimilation of dynamic observation sources. When incorporating a new data source, the system only requires fine-tuning the foundation model to function as the respective observation operator, followed by fine-tuning the corresponding DA model within the proposed cascaded DA framework. Conversely, if an observation source becomes temporarily unavailable, the corresponding model can be excluded from the pipeline without necessitating a full-system retraining. These capabilities are critical for advancing ML-based DA and forecasting toward practical operationalization. This achievement highlights the growing role of AI in advancing operational weather forecasting and establishes XiChen as a foundation for leveraging data-driven approaches to better understand and predict Earth system dynamics with broad societal and environmental relevance. Rather than serving as independent alternatives, ML-based systems such as XiChen are best characterized as sophisticated extensions of existing NWP infrastructure. This complementary relationship stems from their fundamental reliance on NWP-generated reanalysis fields for training, positioning ML models as a vital augmentation of physics-based forecasting.
@@
Although XiChen demonstrates significant potential for practical applications, it also exhibits certain limitations. First, the model experiences smoothing effects during both the DA and weather forecasting processes, a common challenge in ML-based weather forecasting models [bonavita2024some]. Potential solutions include designing more suitable loss functions [xu2024extremecast], incorporating physical constraints [xu2024weathergft], or employing post-processing techniques based on generative models [zhong2024fuxiextreme]. Second, the current implementation of XiChen integrates satellite observations from only four satellite instruments (MHS, ASCAT, AMSU-A, and SATWND), a limited subset compared to the extensive data streams utilized in operational NWP systems [zhong2024fuxiextreme]. In the future, incorporating additional indirect observations could significantly enhance the performance of the analysis field. Third, the model produces only deterministic weather forecasts. However, making critical decisions requires awareness of all possible weather scenarios and their associated probabilities. Therefore, it is necessary to extend the model to a probabilistic version. Addressing this limitation in the future may involve reconstructing the model to adopt a probabilistic framework [price2025probabilistic] or incorporating perturbations into the XiChen system [chen2024machine]. Finally, while XiChen demonstrates impressive performance in DA and medium-range global weather forecasting tasks, its deployment in operational settings would require addressing additional challenges related to real-time data ingestion, model updates, and integration with existing NWP infrastructure.
4 Method